OpenStudy (anonymous):

Simplify quantity 1 over 36 minus 1 over x squared all over 1 over 6 plus 1 over x.

OpenStudy (anonymous):

Can you put into the equation format?

OpenStudy (mathstudent55):

Is this it? \( \dfrac{ \dfrac{1}{36} - \dfrac{1}{x^2} } {\dfrac{1}{6} + \dfrac{1}{x}}\)

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

Step 1: Find the LCD of every small denominator: 36, x^2, 6, and x. Do that. What do you get?

OpenStudy (anonymous):

36x

OpenStudy (anonymous):

6x?

OpenStudy (mathstudent55):

The 36 is correct, but since you have x^2 and x, the LCD is 36x^2.

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (mathstudent55):

Step 2: Multiply the numerator and denominator of the main fraction by the LCD you just found in Step 1.

OpenStudy (mathstudent55):

\( =\dfrac{36x^2}{36x^2} \times \dfrac{ \dfrac{1}{36} - \dfrac{1}{x^2} } {\dfrac{1}{6} + \dfrac{1}{x}}\)

OpenStudy (mathstudent55):

When you multiply fractions together, you multiply the numerators and yo9u multiply the denominators. \( =\dfrac{ 36x^2 \left(\dfrac{1}{36} - \dfrac{1}{x^2}\right) } {36x^2 \left(\dfrac{1}{6} + \dfrac{1}{x} \right)}\)

OpenStudy (mathstudent55):

Now you use the distributive property in the numerator and in the denominator. \( =\dfrac{ 36x^2 \times \dfrac{1}{36} - 36x^2 \times\dfrac{1}{x^2} } {36x^2 \times \dfrac{1}{6} + 36x^2 \times\dfrac{1}{x} }\) Now all small denominators cancel out. \( =\dfrac{ \cancel{36}x^2 \times \dfrac{1}{\cancel{36}} - 36\cancel{x^2} \times\dfrac{1}{\cancel{x^2}} } {\cancel{36} 6 x^2 \times \dfrac{1}{\cancel{6}} + 36\cancel{x^2}x \times\dfrac{1}{\cancel{x}} }\) After all this you get you get: \(=\dfrac{x^2 -36}{6x^2 + 36x} \) Now we factor the numerator and denominator: \(=\dfrac{(x + 6)(x - 6)}{6x(x + 6)} \) Finally we reduce this fraction to get: \(=\dfrac{\cancel{(x + 6)}(x - 6)}{6x\cancel{(x + 6)}} \) \(\dfrac{x - 6}{6x} \)